b) Is the cost of the solar project justified by the savings? (10 marks) QUESTION 2 (25 MARKS) For several years, CC Corporation has purchased the carafe assembly of its major coffeemaker line at an annual cost of RM 11 million. The suggestion to make the component in-house has been made. For the three departments involved, the annual indirect cost rates, estimated material, labor and hours are found in Table Q2. The allocated hours column is the time necessary to produce the carafes for a year. Equipment must be purchased with the following estimates: first cost of RM 10 million, salvage value of RM 250,000, and life of 10 years. Perform an economic analysis for the make alternative, assuming that a market rate of 15% per year is the MARR. (25 Marks)
Answer #1

Answer:

Given

One time cost A= RM 10 million=RM 10,000,000

Salvage Value S =RM 250,000

Project Life n= 10 years

Annual Cost P= RM 11 miilion=RM 11,000,000

MARR r=15% per year

PV = -PV of Annual cost+PV of salvage cost- One time cost

PV of Annual cost= P*(1-(1+r)^-n)/r=11,000,000*(1-(1+15%)^-10)/15%=RM 55,206,454.88

PV of salvage value= S/(1+r)^n=250000/(1+15%)^10=RM 61796.18

PV =-55,206,454.88+61796.18-10,000,000=-RM 65,144,658.71

For make in house machine PV of in house machine   should be less than PV of purchased machine.